OpenStudy (samigupta8):

Lim{1+(arc cosx)^1-x } when x tends to 1-

OpenStudy (perl):

please show any work done so far

OpenStudy (samigupta8):

Lt (1+(arc cosx)^h)

OpenStudy (samigupta8):

I replaced x by 1-h

OpenStudy (samigupta8):

Sry (1+(arc cos(1-h))^h)

OpenStudy (perl):

Is this the question $$ \Large \lim_{x \to 1^{-}} 1 + (\arccos x)^{1-x} $$

OpenStudy (samigupta8):

Yaa correct

OpenStudy (samigupta8):

After this i m not getting d way out to solve this...

OpenStudy (perl):

where does h go, as x -> 1-

OpenStudy (samigupta8):

We can write x as 1-h

OpenStudy (perl):

where does h tend to as x approaches 1 from below (or from the left of 1)

OpenStudy (samigupta8):

h tends to 0

OpenStudy (samigupta8):

i have assumed that h is very minimal term as that we could replace x by 1-h

OpenStudy (perl):

yes you can do that.

OpenStudy (samigupta8):

But after that how to move on with this

OpenStudy (perl):

I don't see how we can move on at this point. have you thought about using a logarithm

OpenStudy (samigupta8):

Well ,for the term (arc cos(1-h))^h we can use that

OpenStudy (perl):

For instance, we use logs to find the limit of $$ \large\lim_{x \to 0^{+}} x^x $$

OpenStudy (samigupta8):

But the biggest fact is that i really didn't think about it unless u have asked me to think on it

OpenStudy (perl):

if you plug in h = 0 into (arc cos(1-h))^h we get (arc cos(1-0))^0 = ( arcos 1 ) ^0 = 0 ^0 which is indeterminate form

OpenStudy (samigupta8):

In such forms only we use logarithm 0^0 or infinity ^ infinity

OpenStudy (perl):

correct, thats a standard approach for limit x->a f(x)^g(x) , as f(x) ->0 , g(x) ->0 or you can find the limit of e^ln(f(x)^g(x))

OpenStudy (perl):

would you like to use logs on your expression?

OpenStudy (samigupta8):

So we would be left with log y= h log arc cos(1-h)

OpenStudy (samigupta8):

I don't know how to proceed frm here

hartnn (hartnn):

by taking logs on both sides, we convert a 0^0 expression into 0*0 expression, wherein we can now plug in the limiting value in the function :)

hartnn (hartnn):

in an easy language, now just put h=1 :P

OpenStudy (samigupta8):

What is P here?

OpenStudy (samigupta8):

Oh sorry i don't get it at frst

OpenStudy (samigupta8):

Bt h tends to 0 how can we put 1 now

hartnn (hartnn):

sorry, typo then. plug in whatever h approaches to :)

OpenStudy (samigupta8):

Log 0 is not defined then

hartnn (hartnn):

as \(x \to 1^- \\ h \to 0^+\) do you understand this?

OpenStudy (samigupta8):

Yup

OpenStudy (samigupta8):

So u want to say that 1-h is not exactly 1 it is sumthing 0.99999 sort of and its cos inverse will lead me to sumthin 0.00001 sort

hartnn (hartnn):

yes, but still log of that is a very high negative number \(\to -\infty \)

OpenStudy (samigupta8):

Yup so what now....

OpenStudy (samigupta8):

U said that it would be 0

OpenStudy (samigupta8):

Guys what now........

OpenStudy (perl):

you can apply Lhopitals rule to an expression 0 * (-infinity)

OpenStudy (samigupta8):

The formula can be used bt this would be complicated as again we r going to be left vid 0/0 form

OpenStudy (perl):

$$ \lim_{h \to 0^{+}} ~ 1 + (\arccos(1-h))^h \\ \ \\ = \lim_{h \to 0^{+}} ~ 1 + \lim_{h \to 0^{+}} (\arccos(1-h))^h \\ \ \\ = 1 + \lim_{h \to 0^{+}} (\arccos(1-h))^h $$

OpenStudy (perl):

we only need to concentrate on finding the limit of \( \lim_{h \to 0^{+}} (\arccos(1-h))^h \)

OpenStudy (samigupta8):

Yaa correct

OpenStudy (samigupta8):

This is what i n @hartnn discussed...:)

OpenStudy (samigupta8):

I got upto this point

OpenStudy (perl):

suppose that $$ \lim_{h \to 0^{+}} (\arccos(1-h))^h = y \\ \ \\ \text { then } \\ \ \\ \ln (y) = \ln( \lim_{h \to 0^{+}} (\arccos(1-h))^h) \\ \ \\ =\lim_{h \to 0^{+}} \ln (\arccos(1-h))^h \\ \ \\ =\lim_{h \to 0^{+}} h \cdot \ln (\arccos(1-h))= 0 \cdot (-\infty) \\ \ \\ =\lim_{h \to 0^{+}} \frac{ \ln (\arccos(1-h))}{1 / h } \\ \ \\ \implies $$

OpenStudy (perl):

now you can use Lhopitals rule

OpenStudy (samigupta8):

Yaa right....this last step was also apt...bt when we wil apply l hospital then again 0/0 form is leftover

OpenStudy (perl):

it shouldn't if you did it correctly

OpenStudy (perl):

this is actually more difficult than it needs to be because you made that substitution earlier it would be easier with the original x variable

OpenStudy (samigupta8):

That's what my professor used to tell me that it"s bettr to convert into h .This is why i started with all that...

OpenStudy (perl):

not in every case is it better to convert to h in some cases it does make it simpler, in others it does not

OpenStudy (samigupta8):

Yaa correct i have got dis...

OpenStudy (samigupta8):

Now when we will substitute h as 0 it would give us 0/0 form

OpenStudy (perl):

lets simplify first

OpenStudy (perl):

did you get this? yes you are right , this is another indeterminate form. we can apply Lhopital again, if that doesn't work, then drop it and try a different approach.

OpenStudy (samigupta8):

I told u already.....:)..anyways u r right we can try that

OpenStudy (samigupta8):

U put 1-(1-h)^2 as h^2 ...How?

OpenStudy (perl):

thats a major typo

OpenStudy (samigupta8):

??

OpenStudy (perl):

\[\text{if} ~~ \lim_{h \to 0^{+}} (\arccos(1-h))^h = y \\ \ \\ \text { then } \\ \ \\ \ln (y) = \ln( \lim_{h \to 0^{+}} (\arccos(1-h))^h) \\ \ \\ =\lim_{h \to 0^{+}} \ln (\arccos(1-h))^h \\ \ \\ =\lim_{h \to 0^{+}} h \cdot \ln (\arccos(1-h))= 0 \cdot (-\infty) \\ \ \\ =\lim_{h \to 0^{+}} \frac{ \ln (\arccos(1-h))}{1 / h } \\ \ \\\Large \implies \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1- (1-h)^2} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1- (1 - 2h + h^2)} } \cdot (-1)}{-1 / h^2 } \]

OpenStudy (samigupta8):

Yes now its fine

OpenStudy (samigupta8):

Now actually it would be h^2 whole upon arc cos(1-h)√2h-h^2

OpenStudy (perl):

$$\Large \implies \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1- (1-h)^2} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1-1 + 2h -h^2)} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ 2h -h^2} }}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ h( 2 -h)} }}{-1 / h^2 } $$

OpenStudy (samigupta8):

We can take h^2 common from root and we will be left with h√2/h-1

OpenStudy (perl):

\[\Large \implies \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1- (1-h)^2} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1-1 + 2h -h^2)} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ 2h -h^2} }}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ h( 2 -h)} }}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ h( 2 -h)} } \cdot \frac{-h^2}{ 1 } \]

OpenStudy (samigupta8):

Hey...hey....it's coming out to be 1 if we cancel h^2 with the h in denominator n then 2/h would be infinite and this term is in denominator we will get it to be 0 now log 1 is 0 so its limit is 1

OpenStudy (perl):

\[\Large \implies \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1- (1-h)^2} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1-1 + 2h -h^2)} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ 2h -h^2} }}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ h( 2 -h)} }}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{1}{\arccos(1-h) } \cdot \frac{1}{\sqrt{ h} \sqrt{ 2 -h} } \cdot \frac{-h^2}{ 1 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{-h^{3/2}}{\arccos(1-h) \sqrt{2-h} } \]

OpenStudy (perl):

can you draw that out?

OpenStudy (samigupta8):

I m not getting how to draw it out....

OpenStudy (perl):

The draw tool?

OpenStudy (samigupta8):

But i m not sure if we can take h2 common in the root or not...

OpenStudy (samigupta8):

Bcz my tchr has tolde we can do so when limit tends to infinite then we take out terms with higher power common n solve the limit thereafter...

OpenStudy (samigupta8):

What say @hartnn ?

OpenStudy (perl):

the limit should come out to be zero since ln(y) = 0 implies y = e^0 = 1

OpenStudy (samigupta8):

Bt am i doing it right i mean can we take common like that...i m confused over it

OpenStudy (perl):

I don't see how you can do that. can you explain it again

OpenStudy (perl):

If we go back a few steps, we can try to introduce 1/h^2 inside the radical.

OpenStudy (samigupta8):

Yes you try to do that take h^2 common in the radical itself and solve it

OpenStudy (perl):

\[\Large \implies \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1- (1-h)^2} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ \frac{1}{\arccos(1-h) } \cdot \frac{-1}{\sqrt{ 1-1 + 2h -h^2)} } \cdot (-1)}{-1 / h^2 } \\ \ \\ \Large = \lim_{h \to 0^{+}} \frac{ -h^2 }{\arccos(1-h) {\sqrt{ 2h -h^2} }} \\ \ \\ \]

OpenStudy (perl):

that would still leave an h in the numerator

OpenStudy (perl):

I don't think the substitution made was a productive substitution.

OpenStudy (perl):

There are two things you can do 1. go back and use the original x, dont make substitution 2. make a linear (or quadratic) approximation for arccos(x) near x = 1

OpenStudy (samigupta8):

Linear approximation in what sense ....

OpenStudy (perl):

for example if you want to find $$\large \lim_{x \to 0^{+}} (\sin x )^ x \approx \lim_{x \to 0^{+}} (x )^ x $$ you can approximate sin x by x

OpenStudy (perl):

y=sin x behaves like y=x , as x approaches zero

OpenStudy (samigupta8):

Bt in this particular case what sort of approximation are you gonna use...

OpenStudy (perl):

what function does arccos x behave like, as x approaches 1

OpenStudy (perl):

im going to try this problem using the original variable x instead of h

OpenStudy (samigupta8):

Actually u r the first person who has told me about such a method so i really don't know such a function....

OpenStudy (samigupta8):

:) i m really sorry but i m taking a whole lot of ur precious time in this one question

OpenStudy (samigupta8):

Alryt u do it ur way ....

OpenStudy (perl):

\[ \Large \lim_{x \to 1^{-}} \ln(\arccos x) ^{1-x} = \lim_{x \to 1^{-}} (1-x) \cdot \ln(\arccos x) \\ \ \\ \Large = \lim_{x \to 1^{-}} \frac{\ln(\arccos x) }{\frac{1}{1-x}} \\ \ \\ \Large \implies \lim_{x \to 1^{-}} \frac{ \frac{ -1 }{\arccos x~ \sqrt{1-x^2} } } { \frac{1}{(1-x)^2} } \]

OpenStudy (samigupta8):

This is also 0/0 form

hartnn (hartnn):

would it be easier if we bring arccos in the denominator? (1-x) divided by 1/ln (arccos x)

OpenStudy (perl):

\[ \large \lim_{x \to 1^{-}} \ln(\arccos x) ^{1-x} = \lim_{x \to 1^{-}} (1-x) \cdot \ln(\arccos x) \\ \ \\ \large = \lim_{x \to 1^{-}} \frac{\ln(\arccos x) }{\frac{1}{1-x}} \\ \ \\ \large \implies \lim_{x \to 1^{-}} \frac{ \frac{ -1 }{\arccos x~ \sqrt{1-x^2} } } { \frac{1}{(1-x)^2} } = \large \lim_{x \to 1^{-}} \frac{ -1 }{\arccos x~ \sqrt{1-x^2} } \cdot \frac{(1-x)^2}{1} \]

OpenStudy (perl):

hartnn it might be, didn't try that :)

OpenStudy (samigupta8):

Isn't this one question just too much for each of us...